# Optimizing Multi-level Magic State Factories for Fault-Tolerant Quantum Architectures

### v1

@@ IV.2.1 Set the Core Area’s Code Distance
IV.2.2 Set the First-Level’s Code Distance
Once $d_{L+1}$ has been determined, we calculate $E_{\text{core}}$ using Eq. 6 and, following Eqs. 1 and 11, we set the magic state error budget according to
@@ IV.2.2 Set the First-Level’s Code Distance
Here, the objective function (38) is a function for the magic state output error rate as in Eqs. 14 and 15, and the constraint (39), determined for example by Eq. 28, ensures that the acceptance probability is positive in order to validate the choice. While this problem can be solved using gradient-based methods, the limited solution space affords a quick way of verifying optimality by iterating through code distances while verifying feasibility until the turning point is reached in both functions $e_{\text{out},1}$ and $P$.
IV.2.3 Determine the Number of Distillation Levels
Minimizing the number of distillation levels $L$ generally leads to space-efficient solutions by avoiding the overhead of added levels even if they have small code distances. We decide on the number of distillation levels by incrementally adding levels with a code distance of $d_{\ell}=\infty$, which results from the output error bound from the distillation protocol, until the magic state error rate $e_{\text{out},L}\leq e_{\text{msf}}$, considering the accumulated errors described in Section III.2.
IV.2.4 Update the Code Distances for All Levels
@@
IV.2.5 Determine the Number of Distillation Units
@@
$u_{L}=\left\lceil\dfrac{\beta D_{L}}{C_{L+1}}\right\rceil.$ (41)
@@
This process is repeated for each level, with $\beta$ adjusted based on the production rate of the selected number of units. Thus, when the chosen number of distillation units $u_{\ell}$ leads to a production rate above the target $\beta$, $u_{\ell-1}$ takes into consideration this gap and selects only enough units to approximate the target slowdown rate.
IV.2.6 Local Search
To further improve the solution, a local search is conducted. The algorithm increases the code distance one level at a time, rebalancing production and consumption rates while recalculating space costs. If an improvement is found, the solution is updated, and the process continues. The search stops when no further local improvements can be found.
V Experiments
@@
We conducted a set of experiments to analyze optimal solutions for the studied problem and to perform a sensitivity analysis on the parameters that were input, including the number of logical qubits ($Q$), the circuit depth ($T$), and hardware noise parameters. For simplicity we assumed $e_{\text{cliff}}=e_{\text{mem}}$ and that the quantum memory protocol had $r=d$ rounds everywhere, therefore, the hardware noise parameters were merely the quantum memory error suppression rate ($\Lambda_{\text{mem}}$), and the preparation error rate ($e_{\text{prep}}$). The experiments did not require much computational power; they should be easily reproducible even on a personal computer. We also assumed the 15:1 distillation protocol with the layout of Fig. 3 were used for all $\forall\ell\in\{1,\ldots,L\}$; therefore, $M=15$, $N=1$, $P=1-15e_{\text{in}}-356e_{\text{mem}}$, $Q=13$, and $e_{\text{out}}=35e_{\text{in}}^{3}+7.1e_{\text{mem}}$.
@@
As part of the default settings, we considered the error budget $E=1\%$. A $\Lambda_{\text{mem}}=10$ model was simulated using 1QBit’s TopQAD software [27], resulting in $\mu_{\text{mem}}=3.8\times 10^{-3}$ for the logical error rate model, and $\mu_{\text{prep}}=1.44\times 10^{-4}$ and $\Lambda_{\text{prep}}=-2.5\times 10^{-5}$ for the magic state preparation model Eq. 20. Finally, the parity-check time was set to $W=450$ nanoseconds. All experiments were conducted with $\alpha=0.1$, as preliminary results showed that variations in $\alpha$ had minimal impact on the solution to the problem.
V.1 Space and Time Cost Estimates
Figure 6 presents the space and time cost estimates for a quantum circuit with one billion non-Clifford gates and 1000 logical data qubits. Each point in the figure represents a solution to the architecture design problem for a predefined value of $\beta$, ranging from $\beta=0.2$ to the case where each distillation level has only one unit. The Pareto frontier is highlighted, showing the optimal trade-offs between space and time costs. The solutions on this frontier range from 1.9 seconds and 4.5 million physical qubits (for $\beta=0.2$) to 1.7 minutes and 2.7 million physical qubits (for $\beta\approx 10$).
@@ V.1 Space and Time Cost Estimates
Figure 6: Space and time cost estimates for a circuit with $T=10^{6}$ and $Q=10^{3}$ and computers with $\Lambda_{\text{mem}}=10$. The Pareto front points are highlighted, indicating the optimal solutions for the architecture design problem. The point where $\beta=1$ represents the serial scheduling case. The points to the left of $\beta=1$ assume different degrees of circuit parallelization, while those to the right reflect increasing $\beta$ values until all distillation levels have only one unit.
@@
In the best-case time scenario, assuming a highly parallelizable circuit (for $T_{\text{depth}}=0.2T$), a 5×5$\times$5 × speedup could be achieved with 50% more qubits compared to serial execution (for $\beta=1$). The trend in the plot demonstrates that physical qubit counts increase rapidly as runtime is further reduced, indicating how costly it is to implement the extra resources needed in an MSF to produce magic states sufficiently quickly to allow performing operations in the core processor in parallel whenever possible.
@@
Looking in the other direction, the space-optimal solution reduces the number of physical qubits by 11% compared to serial scheduling, and it slows down execution by a factor of 10×10$\times$10 ×. Both solutions require two distillation levels (i.e., $L=2$), but, for $\beta=1$, the number of distillation units is $u\in\{56,9\}$, whereas for $\beta=10$, it is reduced to $u\in\{6,1\}$. This represents a significant reduction in MSF size—from around 386,000 to 39,000 physical qubits—given the code distances $d\in\{7,17\}$. The two spikes observed for the physical qubit estimates in the plot of Fig. 6 reflect increases of $d_{L+1}$ from 23 to 25 and then to 27.
V.2 Circuit Parameter Sensitivity
@@
Interestingly, when analyzing circuits of different sizes, we observe that the space requirements for running relatively smaller or larger circuits are comparable. For example, for $Q=10^{2}$, the space cost for $T=10^{6}$ ranges between $10^{5}$ and $10^{6}$ physical qubits, while for $T=10^{15}$, the space requirements increase by less than an order of magnitude. In the cases where serial scheduling is employed, physical space requirements increase from approximately $7\times 10^{5}$ to $2\times 10^{6}$ qubits.
V.3 Hardware Noise Parameter Sensitivity
@@
Next, we performed experiments to analyze the sensitivity of space and time estimates to changes in the hardware noise profile, particularly, by varying the error suppression factor $\Lambda_{\text{mem}}$ and magic state preparation error rate $e_{\text{prep}}$. Figure 1(b) shows the results for a circuit with $T=10^{12}$ and $Q=100$, considering the combinations of $\Lambda_{\text{mem}}\in\{3,5,10\}$ and preset values of the magic state preparation error rates $e_{\text{prep}}\in\{10^{-2},10^{-5},10^{-8}\}$.
@@
The figure shows that, under a scenario where the circuit has a parallelization potential leading to a 5×5$\times$5 × speedup compared to the serial case, the circuit could be run in 1–2 months depending on the quality of the physical qubits and the prepared magic states. The primary bottleneck in terms of time is the increased code distance in the core processor due to lower-quality qubits. For example, the worst-case scenario requires a 2.5 times higher encoding for the core processor than the best-case scenario. This difference also affects space requirements—5–10 times fewer physical qubits are needed in the case where the error suppression factor and magic state preparation error rate are of their best values as compared to the worst.
This analysis also shows the trade-offs between improving magic state preparation error rates and enhancing the error suppression factor. For instance, with a one-year runtime budget, improving magic state preparation reduces the physical qubit requirement from $2\times 10^{7}$ to $1\times 10^{7}$, while improving the suppression factor yields a greater reduction, that is, to $1\times 10^{6}$ for the scenarios tested.
VI Conclusion
@@
The main takeaway from our experiments is that quantum resource estimates can be quickly generated when a few key parameters are known, such as the circuit volume (in terms of $\pi/8$ rotation gates and logical qubits requested), the error suppression rates of the quantum error correction scheme, and the specific protocols used for magic state preparation, growth, and distillation. Although certain simplifying assumptions were made to facilitate model development, such as using the same distillation protocols across all distillation levels, the proposed framework is adaptable to other quantum architectures, including alternative MSF designs and other resource state factories.
We observed that, depending on the circuit and hardware characteristics, the quantum error correction scheme in the proposed architecture requires between 100,000 and 100 million physical qubits. Lower qubit counts (as low as 10,000) are achievable in scenarios with near-perfect magic state preparation or very high quality QPUs. However, decade-long execution times are still prohibitive for utility-era quantum computers. To mitigate this issue, future research should focus on discovering novel quantum algorithms, improving circuit synthesis and compilation, and drastically improving hardware quality.

### v2

@@ III.2.1 Set the Core Area’s Code Distance
III.2.2 Set the First-Level’s Code Distance
Once $d_{L+1}$ has been determined, we calculate $E_{\text{core}}$ using Eq. 29 and, following Eqs. 1 and 34, we set the magic state error budget according to
@@ III.2.2 Set the First-Level’s Code Distance
Here, the objective function (21) is a function for the magic state output error rate as in Eqs. 37 and 38, and the constraint (22), determined for example by Eq. 8, ensures that the success probability is positive in order to validate the choice. While this problem can be solved using gradient-based methods, the limited solution space affords a quick way of verifying optimality by iterating through code distances while verifying feasibility until the turning point is reached in both functions $e_{\text{out},1}$ and $P$.
III.2.3 Determine the Number of Distillation Levels
Minimizing the number of distillation levels $L$ generally results in more space-efficient solutions by avoiding the overhead associated with additional levels, even those with small code distances. The magic states produced by the MSF must achieve an error rate below the threshold defined in Eq. 34, that is, $e_{\text{out},L}\leq e_{\text{msf}}$. Considering the accumulated errors described in Section B.2, the error rate of the output magic states after a distillation round is bounded by $35e_{\text{in},\ell}^{3}$ for the 15:1 distillation protocol under the assumption that $d_{\ell}\rightarrow\infty$, since $e_{\text{cliff},\ell}\rightarrow 0$ as derived from Eqs. 32 and 38. Consequently, to minimize $L$, it suffices to incrementally add distillation levels with $d_{\ell}=\infty$ until the magic state output error rate satisfies the required threshold.
III.2.4 Update the Code Distances for All Levels
@@
III.2.5 Determine the Number of Distillation Units
@@
$u_{L}=\left\lceil\dfrac{D_{L}}{\beta C_{L+1}}\right\rceil.$ (24)
@@
This process is repeated for each level using Eq. 7 for the consumption rate of each level in the MSF with $\beta$ adjusted based on the production rate of the selected number of units. Thus, when the chosen number of distillation units $u_{\ell}$ leads to a production rate above the target $\beta$, $u_{\ell-1}$ takes into consideration this gap and selects only enough units to approximate the target slowdown rate.
III.2.6 Local Search
To search for further improvements to the solution, a local search is conducted by incrementally increasing the code distance at each level, one level at a time, rebalancing production and consumption rates, and recalculating space costs. If an improvement is found, the solution is updated, and the process continues. The search stops when no further local improvements are possible. This process enables our optimization framework to always find optimal solutions for cases where $L\leq 2$, which correspond to the most common architectures in the experiments we conducted. For $L\geq 3$, the local search can be replaced with a greedy search that checks all possible combinations of code distances, ensuring optimality, if desired, but requiring additional classical computation.
IV Numerically Studied Space–Time Trade-offs
@@
We conducted a set of experiments to analyze optimal solutions for the studied problem and to perform a sensitivity analysis on the parameters that were input, including the number of logical qubits ($Q$), the circuit depth ($T$), and hardware noise parameters. For simplicity we assumed $e_{\text{cliff}}=e_{\text{mem}}$ and that the quantum memory protocol had $r=d$ rounds everywhere, therefore, the hardware noise parameters were merely the quantum memory error suppression rate ($\Lambda_{\text{mem}}$), and the preparation error rate ($e_{\text{prep}}$). The experiments did not require much computational power; they should be easily reproducible even on a personal computer. We also assumed the 15:1 distillation protocol with the layout of Fig. 2 were used for all $\forall\ell\in\{1,\ldots,L\}$; therefore, $M=15$, $N=1$, $O=13$, $P=1-15e_{\text{in}}-356e_{\text{mem}}$, and $e_{\text{out}}=35e_{\text{in}}^{3}+7.1e_{\text{mem}}$.
@@
As part of the default settings, we considered the error budget $E=1\%$. We use the “Target” transmon qubit array parameters presented in Ref. [18] as an example (see their Table 1 for $T_{1}$ and $T_{2}$ times, and gate, preparation, measurement, and reset times, and their respective error rates). We perform circuit-level noise model simulations of rotated surface codes with these specifications using 1QBit’s TopQAD [19] and obtain $\Lambda_{\text{mem}}=9.3$ and $\mu_{\text{mem}}=1.9\times 10^{-2}$ and a magic state preparation error rate of $e_{\text{prep}}=4.73\times 10^{-5}$. The parity-check circuits for this hardware take $W=350$ nanoseconds and we set the reaction time to $\gamma=10\ \mu s$. All experiments were conducted with $\alpha=0.1$, as preliminary results showed that variations in $\alpha$ only had negligible impacts on the results.
IV.1 Space and Time Cost Estimates
@@
Figure 3: Space and time cost estimates for varying magic state factory sizes. Estimates for the FeMoco-76 circuit with $T=1.4\times 10^{13}$ and $Q=1972$m and quantum computers with $\Lambda_{\text{mem}}=9.3$. The Pareto frontier points are highlighted, indicating the optimal solutions for the assembly problem. The point where $\beta=1$ represents the serial scheduling case. The points to the left of $\beta=1$ assume different degrees of circuit parallelization, while those to the right reflect increasing $\beta$ values until all distillation levels have only one unit. The two sharp increases in the physical qubit counts are due to increases in code distances in the core processor needed to meet the error budget at extended idling times.
We provide space and time cost estimates for electronic-structure quantum computations for the 76-orbital FeMoco molecule (FeMoco-76) with a precision in energy estimation of 0.1 mHa generated using Double Factorization Qubitization, as presented in Ref. [18]. The resulting quantum circuit is estimated to require the execution of $1.4\times 10^{13}$ $T$ gates using 1972 logical data qubits. This circuit serves as a case study to demonstrate the capabilities of our methods. It is important to note that electronic-structure quantum computation is an active field of research, and circuits with reduced logical resource requirements may be achievable using alternative algorithms or by relaxing chemical accuracy constraints [20].
Figure 3 presents the space and time cost estimates for the FeMoco-76 circuit. Each point in the figure represents a solution to the assembly problem for a predefined value of $\beta$, ranging from $\beta=0.2$ to the case where each distillation level has only one unit. The Pareto frontier is highlighted, showing the optimal trade-offs between space and time costs. The solutions on this frontier range from 1.2 years and 27.2 million physical qubits (for $\beta=0.2$) to 12.4 years and 15.1 million physical qubits (for $\beta\approx 2$).
@@
In the best-case time scenario, assuming a highly parallelizable circuit (for $T_{\text{depth}}=0.2T$), a 5×5$\times$5 × speedup could be achieved with 66% more qubits compared to serial execution (for $\beta=1$). The trend in the plot demonstrates that physical qubit counts increase rapidly as runtime is further reduced, indicating how costly it is to implement the extra resources needed in an MSF to produce magic states sufficiently quickly to allow performing operations in the core processor in parallel whenever possible.
@@
Looking in the other direction, the space-optimal solution reduces the number of physical qubits by 8.2% compared to serial scheduling, and it slows down execution by a factor of 2×2$\times$2 ×. Both solutions require two distillation levels (i.e., $L=2$), but, for $\beta=1$, the number of distillation units is $u\in\{72,14\}$, whereas for $\beta=2$, it is reduced to $u\in\{36,7\}$. This represents a significant reduction in MSF size—from around 2.7 million to 1.3 million physical qubits—given the code distances $d\in\{15,37\}$. The two spikes observed for the physical qubit estimates in the plot of Fig. 3 reflect increases of $d_{L+1}$ from 41 to 43 and then to 45.
IV.2 Circuit Parameter Sensitivity
(a) Space–time trade-offs with varying circuit parameters.
(b) Space–time trade-offs with varying hardware parameters.
Figure 4: (a) Pareto frontier estimates are shown for combinations of $T$-count $T=10^{6}$ to $10^{15}$ and logical data qubits $Q=10^{2}$ to $10^{4}$ in a quantum circuit, based on an FTQC model with a quantum memory error prefactor of $\mu_{\text{mem}}=1.9\times 10^{-2}$, an error suppression rate of $\Lambda_{\text{mem}}=9.3$, and a magic state preparation error rate of $e_{\text{prep}}=4.73\times 10^{-5}$. While runtime is heavily influenced by the number of non-Clifford gates in the circuit and the MSF size, these estimates demonstrate that quantum computers require between $10^{5}$ and $10^{8}$ physical qubits across all scenarios evaluated. (b) Pareto frontier estimates for combinations of $\Lambda_{\text{mem}}=3$ to $10$ and a magic state preparation error rate of $e_{\text{prep}}=10^{-2}$ to $10^{-8}$ for the FeMoco-76 circuit with $T=1.4\times 10^{13}$ and $Q=1972$. The estimates show that improving $\Lambda_{\text{mem}}$ among the scenarios tested can reduce the physical qubit count by up to one order of magnitude with a smaller impact on the expected runtime due to the reduced code distances. Meanwhile, improving $e_{\text{prep}}$ has a small effect on reducing physical qubit count because of the relatively large size of the core processor compared to the MSF, and no effect on the expected runtime.
@@ IV.2 Circuit Parameter Sensitivity
Interestingly, when analyzing circuits of different sizes, we observe that the space requirements for running relatively smaller or larger circuits are comparable. For example, for $Q=10^{2}$, the space cost for $T=10^{6}$ ranges between $10^{5}$ and $10^{6}$ physical qubits, while for $T=10^{15}$, the space requirements increase by less than an order of magnitude. In the cases where serial scheduling is employed, physical space requirements increase from approximately $6\times 10^{5}$ to $4\times 10^{6}$ qubits.
IV.3 Hardware Noise Parameter Sensitivity
@@
Next, we performed experiments to analyze the sensitivity of space and time estimates to changes in the hardware noise profile, particularly, by varying the error suppression factor $\Lambda_{\text{mem}}$ and magic state preparation error rate $e_{\text{prep}}$. Figure 4(b) shows the results for the FeMoco-76 circuit, considering the combinations of $\Lambda_{\text{mem}}\in\{3,5,10\}$ and preset values of the magic state preparation error rates $e_{\text{prep}}\in\{10^{-2},10^{-5},10^{-8}\}$. Only the Pareto frontier points are shown for each circuit size.
@@
The figure shows that, under a scenario where the circuit has a parallelization potential leading to a 5×5$\times$5 × speedup compared to the serial case, a circuit with $T=1.4\times 10^{13}$ $\pi/8$ rotations could be run in 1–3 years depending on the quality of the physical qubits and the prepared magic states. The bottleneck in terms of time is the increased code distance in the core processor due to lower-quality qubits. For example, the worst-case scenario ($\Lambda=3,e_{\text{prep}}=10^{-2}$) requires a 2.2 times higher encoding for the core processor than the best-case scenario ($\Lambda=10,e_{\text{prep}}=10^{-8}$), that is, $d_{\text{core}}=87$ in the former and $d_{\text{core}}=39$ in the latter, both for the $\beta=0.2$ scenario. This difference also affects space requirements—6.8 times fewer physical qubits are needed in the best-case scenario as compared to the worst (13.9 million vs. 94.4 million).
(a) Space costs
(b) Time costs
Figure 5: Space and time cost estimates for varying reaction times. Estimates for the FeMoco-76 circuit with $T=1.4\times 10^{13}$ and $Q=1972$, and a quantum computer with $\Lambda_{\text{mem}}=9.3$, for $\beta=1$. Estimates are generated for reaction times $\gamma$ ranging from 100 ns to 1 s. The plots indicate that, for the hardware specification assumed, reaction-time-limited computation provides speedups when $\gamma\lesssim d_{L+1}W$, which, for the scenario tested with $d_{L+1}=41$ and $W=350$ ns, results in the threshold at $14.35\mu$s. Beyond this threshold, runtime decreases linearly with decreasing reaction time at a linearly increasing space cost. For slower reaction times, $T$-count-limited computation is optimal. However, longer reaction times increase space costs due to the larger correction qubit storage requirements in the buffer and MSF.
IV.4 Reaction Time Sensitivity
We conducted experiments to analyze how space and time estimates vary with reaction time, that is, the time required to classically process measurement outcomes and determine corrective operations in $\pi/8$ rotations. Figure 5 presents the space and time cost estimates obtained, distinguishing between estimates for the $T$-count-limited computation and the reaction-time-limited computation (as described in Appendix A). The threshold where reaction-time-limited computation can provide a speedup is approximately $\gamma=d_{L+1}W$. Estimates for the $T$-count-limited computation converges to the $\beta=1$ estimates in Figure 3 as reaction time $\gamma\to 0$.
In practical quantum systems, large delays in performing quantum error correction can degrade the error suppression rate for quantum memory. At a certain point, slow communication between quantum and classical systems could push quantum error correction above threshold, i.e., $\Lambda_{\text{mem}}<1$, rendering fault-tolerant computation infeasible. The effects of the reaction time on quantum error correction can be emulated numerically and incorporated into the noise models presented by modelling noise as a function of reaction time, that is, $f_{\text{mem}}(d,r,\gamma)$.
V Conclusion
@@
The main takeaway from our experiments is that quantum resource estimates can be quickly generated when a few key parameters are known, such as the circuit volume (in terms of $\pi/8$ rotation gates and logical qubits requested), the error suppression rates of the quantum error correction scheme, the reaction time, and the specific protocols used for magic state preparation, growth, and distillation. Although certain simplifying assumptions were made to facilitate model development, such as using the same distillation protocols across all distillation levels, the proposed framework is adaptable to other quantum architectures, including alternative MSF designs and other resource state factories.
We observed that, depending on the circuit and hardware characteristics, the quantum error correction scheme in the proposed architecture requires between 100,000 and 100 million physical qubits, scaling to higher numbers when reaction time becomes significantly faster than the logical cycle times to provide computational speedups. When comparing our approach to previous resource estimations in the literature, we found consistent physical resource estimates for utility-scale quantum algorithms with runtimes ranging from seconds to days. Future research may focus on discovering novel quantum algorithms, improving circuit synthesis and compilation, and drastically improving hardware quality to further reduce physical requirements.